# Towards Generalized Video Quality Assessment: A Weak-to-Strong Learning Paradigm

### v3

@@ C.3.1 Probability Modeling
First, we describe how to construct the probability distribution for comparative quality assessments. The comparative token set is defined as:
@@
where $\alpha_{k}$ are fixed weights that reflect the comparative levels. Specifically, the weights are defined as:
@@ C.3.2 Score Modeling
Table 5: Performance of weak-to-strong models trained with pseudo-labels from weak models. For comparison, we also report the performance of our model trained directly on the LSVQ dataset.
# of videos 7,182 3,573 1,200 585 1,020 -
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Weak Teachers
MinimalisticVQA(VII) 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
Weak-to-Strong Students
MinimalisticVQA(VII)-labeled 0.855 0.852 0.762 0.795 0.859 0.857 0.771 0.813 0.808 0.821 0.824 0.833
MinimalisticVQA(IX)-labeled 0.879 0.878 0.794 0.826 0.869 0.871 0.786 0.822 0.843 0.846 0.849 0.859
FAST-VQA-labeled 0.871 0.868 0.785 0.819 0.849 0.855 0.798 0.833 0.825 0.834 0.840 0.850
DOVER-labeled 0.877 0.869 0.780 0.813 0.870 0.875 0.792 0.829 0.819 0.831 0.843 0.850
Q-Align-labeled 0.878 0.876 0.794 0.824 0.873 0.880 0.781 0.825 0.833 0.853 0.848 0.859
Supervised Student
LSVQ-labeled 0.881 0.878 0.797 0.834 0.874 0.874 0.797 0.828 0.830 0.838 0.851 0.861
# of videos 600 357 480 1,200 2,926 -
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Weak Teachers
MinimalisticVQA(VII) 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
Weak-to-Strong Students
MinimalisticVQA(VII)-labeled 0.632 0.717 0.718 0.773 0.318 0.386 0.356 0.412 0.604 0.652 0.536 0.593
MinimalisticVQA(IX)-labeled 0.687 0.748 0.763 0.810 0.383 0.461 0.459 0.515 0.638 0.676 0.591 0.639
FAST-VQA-labeled 0.658 0.766 0.752 0.785 0.392 0.422 0.414 0.493 0.585 0.624 0.550 0.604
DOVER-labeled 0.662 0.758 0.752 0.809 0.449 0.482 0.435 0.519 0.574 0.627 0.554 0.617
Q-Align-labeled 0.671 0.738 0.744 0.785 0.437 0.480 0.450 0.525 0.620 0.668 0.581 0.636
Supervised Student
LSVQ-labeled 0.643 0.713 0.713 0.770 0.451 0.490 0.451 0.485 0.619 0.636 0.577 0.608
@@
With this probability matrix, we estimate the final quality score using maximum a posteriori (MAP) (Tsukida et al., 2011) estimation under Thurstone’s Case V model (Thurstone, 2017). This is formulated as the following convex optimization problem:
@@ Appendix D More Details of Experimental Results
D.1 More Details of Weak-to-strong Generalization Effect
Table 5 presents the per-dataset results from the experiments described in Section 3.3. For in-domain benchmarks, the student model achieves performance comparable to its teachers, with slight improvements, demonstrating that our simple knowledge distillation approach effectively transfers quality assessment knowledge from weak to strong models. For OOD benchmarks, the student model shows substantial improvements over its teachers, highlighting a pronounced weak-to-strong generalization effect.

### v4

@@ C.2.1 Probability Modeling
First, we describe how to construct the probability distribution for comparative quality assessments. For hard ranking, the comparative token set is defined as:
$\mathcal{S}=\{s_{k}\}_{k=1}^{2}=\{\textit{worse},\textit{better}\}.$ (21)
For soft ranking, the comparative token set is extended to:
@@
where $\alpha_{k}$ are fixed weights that reflect the comparative levels. Specifically, for hard ranking, the weights are:
$\{\alpha_{k}\}_{k=1}^{2}=\{0,0.5\}.$ (25)
For soft ranking, the weights are defined as:
@@ C.2.2 Score Modeling
With this probability matrix, we estimate the final quality score using maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53]. This is formulated as the following convex optimization problem:
@@ Appendix D More Details of Experimental Results
We also compare the prediction accuracy of our model with several advanced LLMs, including LLaVA-NeXT-Video (7B) [74], mPLUG-Owl3-8B [68], InternVL2.5-8B [59], Qwen2.5-7B-VL [2], and our base model LLaVA-ov-chat [24], as shown in Table 6. Evidently, our model significantly outperforms all LLMs, suggesting that existing LLMs trained on high-level tasks still struggle with low-level visual perception.